OpenStudy (anonymous):

f(x) = 1/x, f(x)-f(a) / x-a

OpenStudy (phi):

\[ \frac{ f(x+a) - f(x)}{(x+a)-x} \]

OpenStudy (phi):

\[\frac{ \frac{1}{x+a} - \frac{1}{x} }{a} \\ \frac{1}{a}\left( \frac{1}{x+a} - \frac{1}{x}\right)\] now use a common denominator of x(x+a) \[ \frac{1}{a}\left( \frac{x}{x(x+a)} - \frac{x+a}{x(x+a)}\right) \\ \frac{1}{a}\left( \frac{x-x-a}{x(x+a)} \right) \\ \frac{1}{\cancel{a}}\left( \frac{-\cancel{a}}{x(x+a)} \right) \\ - \frac{1}{x(x+a)} \] now take the limit, and let a-->0 \[\lim_{a \rightarrow 0} \frac{ -1 }{ x(x+a) }= \frac{ -1 }{ x^2}= -x^{-2}\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."